### v1

# Adaptive behavior with stable synapses

Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters. However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could also be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks. We argue that such a framework can describe the psychometry of context-dependent tasks on humans and other species, solving the incoherence of plasticity timescales. When the context is changed, the network is dynamically reconfigured, and the predicted output undergoes dynamic updates until it aligns with the information embedded in the context. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
Introduction
The study of behavioral adaptation, observed in both humans and animals, has been a longstanding subject of research. The rapidity with which behaviors can change in response to new cues challenges the conventional explanation of synaptic plasticity. This led researchers to explore mechanisms that underlie such flexible adaptations [1].
In-context learning as an emerging property in AI
In-context learning (ICL) is the capacity of a model to adapt its behavior, without weight updates, to solve tasks not encountered during training. Initially, ICL was observed in architectures tailored for few-shot learning [2] or even zero-shot learning [3]. However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [4, 5]. Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[6, 7, 8]. Despite some intriguing explanations [9, 10, 11], the emergence of this property remains not completely understood. In particular, it is not clear how to associate it with in-context learning abilities observed in biological networks. The major contribution of this work is the definition of a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
Limitations of transformers architecture
Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length. This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding. To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs. Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks. Zucchet et al. [14] explore the efficacy of these deep linear RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and facilitating in-context learning.
Biological support for in-context learning
Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers.
Recent findings on dendritic computational properties [15] and on the complexity of pyramidal neurons dynamics [16] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [17, 18, 19, 20]. It has has been proposed that segregation of dendritic input [18] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [20] would support backpropagation in biological neurons. In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network.
Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of stationary and temporal datasets, as well as for reinforcement learning.
This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
Methods
Dynamical adaptation
Consider a generic learning task, in which the response $y$ to a state $x$ and is influenced by a set of parameters $w$. The latter are adjusted in function of an internal error $e$. This can be rewritten in a generic formulation as follows:
$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f)\\ \tau_{y}\,\dot{y}&=Y(y,w,x)\\ \tau_{w}\,\dot{w}&=W(w,x,e)\end{cases}$ (1)
where $f$ is feedback from the environment (e.g. the reward, the target behavior, …). Alternatively to the standard interpretation of $w$ as synaptic weights of a neuronal network, we consider them as the state variables of a dynamical system coupled to the dynamics of $y$, expressed by the activity of an auxiliary network. For this reason, we also refer to them as “virtual” weights.
Dynamical adaptation for linear regression
Consider now the task of learning a linear relationship between input $x$ and output $y$. In the conventional approach, a neural network is initialized with random parameters, essentially making random guesses. The model estimation can be improved through iterative adjustments of the network parameters (e.g. the synaptic weights), following the gradient of errors evaluated on a training dataset $(x_{i}^{data},y_{i}^{data})$. One step of the gradient would update the weights of a linear model $y=wx$, following $\delta w=-\sum_{i}(y-y_{i}^{data})x^{data}_{i}$, where $\eta$ is the learning rate. By the end of the training process, the network aims at accurately predict the output $y^{test}$ for a new input $x^{test}$. An equivalent error reduction can be achieved by updating the target output $y_{i}^{data}$ to $y_{i}^{data}-wx_{i}^{data}=y_{i}^{data}-y_{i}^{train}$ (let’s denote this the latter quantity as $e_{i}$), where $y_{i}^{train}=wx_{i}^{data}$ is the current model estimation of $y_{i}^{data}$. $e_{i}$ can then be updated iteratively, and when it reaches the value zero, it indicates that the interpretation of the dataset is coherent, and the network predictions on new data $x^{test}$ are reliable. We can interpret this dynamics as changing the way $x^{test}$ is processed in function of the virtual weights $w$. Typically, when a network dynamically alters the way it receives an input, it is referred to as an attention mechanism. Following this concept, our goal is to project the dataset $(x_{i}^{data},y_{i}^{data})$ onto the network, influencing the way a new input $x^{test}$ (projected simultaneously) is processed. The information that drives and influences the attention mechanisms, $(x_{i}^{data},y_{i}^{data})$ in this case, is usually referred to as a context. When the input vector is one dimensional, the task can be solved by a network composed of three modules, one evaluating the error $e$, one evaluating the necessary update $\delta w$, and one evaluating the response $y$ to input data (see Fig. 1A). Their dynamics can be explicitly written as follows  [^1] Here we model the $y$ to be an instantaneous function of the variables $w,x$. This is equivalent to considering $Y(y,w,x)=xw-y$ in Equation (1) and taking the limit $\tau_{y}\to 0$. Similar considerations hold for $e$.:
$\begin{cases}e_{i}&=(y^{data}_{i}-x_{i}^{data}w)\\ \tau_{w}\,\dot{w}&=\sum_{i=1}^{M}x_{i}^{data}\,e_{i}\\ y&=x^{test}w\end{cases}$ (2)
with $e_{i}(0)=y_{i}^{data}-w_{0}x_{i}^{data}$, and $w_{0}$ are the initial weights. The conditioning dataset is $(x_{i}^{data},y_{i}^{data})$, with $i=1,...M$, while $(x^{test},y^{test})$ is one element of the test set. In this case one example ($M=1$) is sufficient to determine the response to new inputs. This formulation is equivalent to what derived in [11], with the difference that here many iteration are recursively evaluated over time, allowing the convergence to the gradient descent solution. See Results section for examples of applications. This approach can be extended to multidimensional linear datasets (see Supporting Information).
Dynamical supervised learning for temporal trajectory
The above formulation is well-posed for stationary dataset, however such a framework (the temporal convergence of the error to zero and of the response to the desired value) is tricky for temporal trajectories, which requires a little change of paradigm, to achieve a formulation that holds at each time step.
First of all we introduce a target temporal trajectory $y^{targ}(t)$, and define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset. In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$. We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$.
This together with Eq. (2) allows us to write:
$\begin{cases}e&=(y^{targ}-y)\\ y&=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{x}\,e\end{cases}$ (3)
where we removed the dependence on time $t$ for simplicity. The operations required are nonlinear, and usually are naturally implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights. However, here $\bm{w}$ are not actual weights but rather dynamical variables.
We propose, as possible implementation of this, that such non-linear functions are computed by two neural networks networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{x},e)\simeq W(\bm{x},e)=\bm{x}\,e\end{cases}$ (4)
In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm).
In addition, we introduce a novel concept known as a gain-modulated network. This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]). These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions. More details in the section "Gain Modulated Reservoir Computing". We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents.
The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$. However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments.
We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions. For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$. This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above. The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time. This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
$\begin{cases}\bm{e}&=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{e}\odot\bm{x}\end{cases}$ (5)
This is the dynamics obtained by evaluating the policy gradient with respect to the virtual weights $\bm{w}$. $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones. $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [23, 24]) at time $t$. It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
This formulation holds for a null discount factor [23, 24]). See Supplementary information for a complete derivation and for the description of the general case. The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (6)
Algorithm Distillation
To demonstrate the ability of this architecture to learn in context, we train networks Y and W using an Algorithmic Distillation protocol, as defined in [25]. Namely, we consider a family of environments $\mathcal{A}$ with different reward distributions  [^2] A family of environments is a set of possible environments with the same action-state transitions but different reward distributions.. We then train the aforementioned networks using $N_{s}$ learning histories obtained through policy gradient methods from a subset of these environments, designated as $\mathcal{A}_{\text{ID}}$. The network performance is then tested out of distribution, i.e., on the rest of the environments $\mathcal{A}_{\text{OOD}}:=\mathcal{A}\setminus\mathcal{A}_{\text{ID}}$
Gain modulated reservoir computing (GM-RC)
Reservoir Computing
Reservoir Computing (RC) represents a paradigm in machine learning that provides a model for biologically plausible computation and learning, drawing inspiration from the information processing mechanisms observed in biological neural networks.
In this approach, a random RNN is employed to extract features from a time-dependent signal $\bm{x}(t)$. The dynamics of the RNN describe the evolution of $N$ hidden units $\bm{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
$\tau_{z}\dot{\bm{z}}=\phi\left(J\bm{z}+R\bm{x}\right)-\bm{z}$ (7)
The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([26]). Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively. Subsequently, the features extracted by the network can be utilized to predict a target signal $\bm{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\bm{y}^{targ}(t)-\Theta\bm{z}(t)|^{2}_{2}$ is minimized. The reservoir is then implementing a map
$\bm{y}(t)=\mathrm{RNN}_{\Theta}(\{\bm{x}(s)\}_{s\leq t})$ (8)
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies. 333 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit. In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\bm{x}\right)$. Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights. This architecture is also referred to in the literature as the Extreme Learning Machine [28].
Gain modulated network architecture
Analysing Equations (4) and (6) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters. Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC). This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [22, 29]. Here, we consider an additional input source $\bm{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$. Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope. Consequently, the resulting RNN equation is formulated as:
$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+R\bm{x})\right)-\bm{z}$ (9)
Where $\bm{b}^{\text{ap}}$ is a constant bias vector. Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the effect of the apical inputs on the network. Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation. The choice for the name $\bm{x}^{\text{ap}}$ is inspired by to current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high level information [22].
Results
Figure 1: Dynamical adaptation to changing context: A. Network architecture composed of 3 modules: $e$, $\delta w$ and $y$. The estimation on new data is produced by $y$, and is modulated by the error variable $e$ through $\delta w$. B. Example of the time course for the three variables in a single realization of the task. The values of $y$ converge to the correct values of (black dashed lines). C. This model accounts for the recovery of optimal performances (bottom, observed in humans and monkeys data) when the context suddenly changes, compared to a slower increase of performances achieved through synaptic plasticity.
Dynamical adaptation replaces synaptic plasticity
To demonstrate that suited responses can be provided relying on dynamical adaptation, we first consider a task requesting to learn a linear relationship between input $x^{data}$ and output $y^{data}$. As described in the Methods we consider a three module model (see Fig. 1A) described by the variables, $e$, $\delta w$, and $y$, whose dynamics is defined by Eq. (2), that is equivalent to following a gradient dynamics as respect to the parameters $w_{0}$. This update can be realized by treating $e$ as an input that, conditioned by the dataset $(x^{data},y^{data})$, evolves over time.
When $e\to 0$ the interpretation of the dataset is coherent, and the input-output mapping performed by the network on the new data is reliable. In this case one example is sufficient to determine the response to new inputs.
Fig. 1B shows the dynamics of three variables: the absolute value of $e$, and $\delta w$ as a consequence the error, are reduced over time, while the $y$ converges to the correct $y^{test}$ given the input $x^{test}$ (shown for 10 different new values).
We show the performances of the network whose parameters $w_{0}$ are trained through a gradient based plasticity on the rule $y^{data}=x^{data}$ (see first bar in the plots Fig. 1C), after the rule is switched to $y^{data}=-x^{data}$ (see, the second bar in which performances drop in both panels). In the top panel $w_{0}$ parameters are trained through gradient updates (we refer to Github repository https://github.com/cristianocapone/ABSS for details on parameters and implementation). In the bottom panel weights are not re-trained, but the last example (that is coherent with the new rule) is projected as a context (e.g. $(x^{data},y^{data})=(-1,1)$, starting the trial numbered as 1). This represents the agent understanding that the rule has been switched, and allows rapid and dynamic network reconfiguration and recovery of performances to optimal values, similarly to what observed in monkeys and humans data. An interesting example is the Wisconsin Card Sorting Test, in which participants are required to sort a deck of cards based on changing rules [30]. Initially, participants are given a rule (e.g., choosing by color) , and after a certain number of trials, the rule is changed without warning (e.g., choosing by shape). Participants then need to adapt to the new rule and adjust their strategy accordingly. The quick adaptation observed in this setup [31] reflects the ability to update mental models of the task, and flexibly apply new rules or strategies (similarly to Fig. 1C-bottom). On the other hand in some pathological subject, the adaptation after rule change appears to be slower [31], suggesting a learning process compatible with gradient-based plasticity (Fig. 1C-top).
Figure 2: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of 10 units (described by the vector $\bm{z}(t)$, depicted in pink) is utilized for learning a periodic trajectory, following the prescription of reservoir computing,. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios.
Dynamical adaptation for temporal trajectories
We aim to demonstrate that such an approach can be generalized to non-stationary scenarios.
We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm, we employ an RNN (Fig. 2A, pink), whose dynamics follows Equation (7), to extract temporal features $\bm{z}(t)$. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 2B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN. The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},\bm{w})\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\,e\end{cases}$ (10)
Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.2A green and orange respectively). Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section. These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$. We refer to Supporting Information for further details on the training procedure.
In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN. This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 2B).
In Fig.2B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line.
The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.2C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C).
Figure 3: Dynamical reinforcement learning: Multi Armed Bandits. A. Schematic of network architecture and task. Virtual weighs parameterize the agent’s policy (orange). At each round, an action (red) is sampled from the policy and played. A reward is then sampled from the current environment. A GM-network (green) then predicts the virtual parameter update. B. Comparing in distribution (ID) and out-of-distribution (OOD) expected reward at the 1000-th round for the policy gradient (PG) training source, reward gain-modulated network ($\gamma>0$), and a network without gain modulation $\gamma=0$. C. ID and OOD model performance varying gain modulation strength ($\gamma$). Solid lines indicate the median score (expected reward), computed over 20 independently trained models. The filled area indicates 25-75% confidence interval. D. Same as B but comparing OOD performance for different network sizes (number of hidden units). E. Robustness to hyperparameter variations in architectures with $(\gamma=0.9)$ and without ($\gamma=0.0$) gain modulation. We report the percentage of models that fail to generalize to OOD environments, varying the standard deviation of apical ($\sigma_{a}$) and basal ($\sigma_{b}$) random input weights. F. Regret comparison between the distilled and the original policy gradient algorithm. We report average regret per round in ID and OOD setting, compared with the one achieved by the PG training source.
Multi Armed Bandits
Dynamical adaptation is now investigated within the framework of reinforcement learning. We provide robust evidence supporting the hypothesis that gain modulation represents a fundamental component in implementing in-context learning in biological agents. We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [32]. Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability. During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 3A, green). This serves as an ideal test bed for evaluating the network ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action, as described in the Methods section. We systematically investigate the impact of the gain modulation on out-of-distribution performance, varying the parameter $\gamma$. Notably, when $\gamma=0$ the reward does not modulate the network gain.
We present our experimental findings in Fig. 3. Initially, we observe that integrating reward gain modulation into the network enables the agent to consistently converge to the optimal policy, even in out-of-distribution environments (see Fig.3B). To further explore this phenomenon, we trained multiple models while varying the strength of the modulation parameter ($\gamma$). We observe that performance improves with higher values of $\gamma$, allowing the generalization of learned rules out-of-distribution (Fig. 3C) and showcasing increased robustness to variations in hyperparameters (Fig. 3E). Furthermore, we observe that in the absence of gain modulation, the network is unable to generalize, merely memorizing the rules from the training set. This is evidenced by its persistence in selecting even-numbered arms even when tested in out-of-distribution environments where high rewards are allocated to odd-numbered arms (see Fig. 3E). Comparing the performances of models of different sizes (number $N$ of hidden units), we see that models with more reward gain modulation need significantly fewer neurons to achieve maximum scores in OOD environments (see Fig.3D).
Analysing the regret curves displayed in Fig. 3F we observe that a model with gain modulation learns a more data-efficient algorithm than its source even in OOD environments. On the contrary, a model without gain modulation fails to generalize to OOD environments, not converging to zero regret (see Fig. 3F, right panel).
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach unobservable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). E. Reward plotted against the number of trials, for the policy gradient (PG) algorithm (displayed in blue lines) applied in the dynamical PG scenario on both training positions (depicted in the left panel) and new positions (shown in the right panel). Thin line: single position, thick line: average across different positions. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [33]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility.
The horizontal and vertical positions of the agent are encoded separately using five input units $\bm{x}$ each, employing Gaussian activation functions with equally spaced centers ($0.4$ apart) and a width of $0.4$.
To achieve this task we consider a network architecture composed of two networks. One GM-network, , $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 4A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig. 4A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equations:
$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq\bm{x}\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\end{cases}$ (11)
Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig. 4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig. 4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions). After multiple trials, the agent successfully achieved precise targeting of the circles. Data collected from these experiments were utilized to train our networks to estimate gradients and scalar products, as defined in Eq. (11).
To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig. 4C, left panel) and new test positions (OOD, Fig. 4C, right panel). For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training. The agent’s precision closely matches that of the plastic policy gradient learning rule. We present the reward plotted against the number of trials for both training (Fig. 4E, right panel, pink lines) and test (Fig. 4C, right panel, pink lines) food locations.
We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig. 4B, D, F). Notably, while performances for ID food locations are acceptable (Fig. 4E, left panel, pink lines), those for OOD cases are extremely low (Fig. 4E, right panel, pink lines). This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
Discussion
It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 34], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8]. Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers. One intriguing explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11]. However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals. In this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
When considering a stationary dataset our model dynamically adapt to the context provided by the data, allowing predictions on new data, without synaptic plasticity. Remarkably, when the learning rule shifts abruptly, the network swiftly adapts without retraining, mimicking human cognitive flexibility observed in tasks like the Wisconsin Card Sorting Test. Conversely, when using gradient-based plasticity, the learning speed resembles what observed in certain pathological conditions [31]. When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward. We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates. Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response. We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
In conclusion, our approach provide an explicit framework to induce ICL in biologically plausible network, possibly opening the route to a formal understanding of ICL in biological agents.
Acknowledgments
This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’) to MM. LF is supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.
Source code availability
The source code is available under CC-BY license in the https://github.com/cristianocapone/ABSS public repository.