### v4

# Adaptive Behavior with Stable Synapses

Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated learning theories to synaptic plasticity. However, such rapid changes are not coherent with the timescales of biological synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical reconfiguration of the network involved, without changing its weights. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks, such as in pyramidal neurons thanks to input segregation and dendritic amplification. We propose a constructive approach to induce in-context learning in an architecture composed of recurrent networks with gain modulation, demonstrating abilities inaccessible to standard networks. In particular, we show that, such architecture can dynamically implement standard gradient-based by encoding weight changes in the activity of another network. We argue that, while these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the level of neural circuits. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
Introduction
Adaptive Behavior in Biological Agents
The study of adaptive behavior in biological agents, including humans and animals, has been a longstanding subject of research [1]. Observations of rapid behavioral shifts in response to environmental cues have raised questions about whether traditional mechanisms, such as synaptic plasticity, can fully account for such flexibility.
A broad range of neural and behavioral data indicates the existence of distinct systems for behavioral choice [2, 3]. The conventional view posits that subcortical areas support habitual, stimulus-based responses, enabling reflex-based and rapid reactions. In contrast, higher cortical areas, such as the prefrontal cortex, are associated with reflective, goal-directed behaviour, which relies on cognitive action planning and deliberate evaluation. Critically, goal-directed behavior allows for the flexible construction of value at each decision point through online planning procedures. This mechanism makes agents immediately sensitive to changes in outcome values, enabling them to dynamically adapt their actions to novel or changing circumstances [4]. Studies in network neuroscience have identified dynamic network reconfiguration, characterized by the flexible recruitment and integration of neural circuits, as a fundamental mechanism supporting this adaptability across cognitive[5] and motor domains[6].
Computational models within the predictive coding framework have recently provided insights into the neural mechanisms underlying dynamic adaptation. These models suggest that adaptation can be understood as a form of Bayesian inference, where the brain continuously infers latent contextual states to guide behavior [7]. A prominent example is the COIN model[8, 9], which posits that adaptation arises through two complementary processes: proper learning, involving the creation and updating of memories, and apparent learning, which dynamically adjusts the expression of pre-existing memories based on inferred context. While these frameworks capture key phenomena of adaptive behavior, such as context-dependent single-trial learning and spontaneous recovery, a significant gap remains between these theoretical models and the neural substrates that implement these computations in the brain.
In-context learning as an emerging property in AI
In machine learning, the capacity of a model to adapt its behavior, without weight updates, is referred to as in-context learning (ICL). Initially, ICL was observed in architectures tailored for few-shot learning [10] or even zero-shot learning [11]. However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [12, 13]. These architectures leverage self-attention, a mechanism that dynamically changes the way past inputs are attended, in order to shape the current response. Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[14, 15, 16]. While there are several intriguing explanations [17, 18] for the emergence of ICL in transformers, a more formal understanding was pioneered by Von Oswald and collaborators [19], who demonstrated that a linear self-attention layer can implement gradient descent steps within its forward pass. This intuition was empirically validated by analyzing trained transformers and was formally proven [20].
Despite their success, transformer architectures face notable shortcomings, particularly in their memory requirements, which scale quadratically with sequence length. These requirements make those architectures largely biologically implausible. Moreover, memory demands constrain their scalability and restrict their applicability to tasks involving long sequences, such as full-length document analysis or video understanding. To overcome these challenges, several efficient deep linear RNN architectures have been proposed [21, 22], demonstrating substantial performance gains over transformers, especially in long-sequence tasks.
However, these deep learning architectures rely on several mechanisms that do not have an obvious correspondence in biological neural systems (such as the attention mechanism itself), leaving an unresolved gap in relating them to the in-context learning abilities exhibited by biological networks. This work bridges this gap by introducing a constructive method to induce in-context learning in biologically plausible neural networks, demonstrating its applicability in a wide range of scenarios.
Biological support for in-context learning
Building on this premise, our investigation explores how in-context learning, as demonstrated in deep learning architectures, might also be realized in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We propose that in-context learning can be facilitated by input segregation and dendritic amplification, two features widely observed in biological networks. We further argue that these mechanisms provide a biologically plausible foundation for implementing a process analogous to the attention mechanism in transformers. Recent findings on dendritic computational properties [23] and on the complexity of pyramidal neurons dynamics [24] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [25, 26, 27, 28]. It has been proposed that segregation of dendritic input [26] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [28] would support backpropagation in biological neurons. Recently, it has been suggested [29] that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network. Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations. This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
Main / Framework description
Dynamical adaptation replaces synaptic plasticity: A popular example
Consider a task inspired by the Wisconsin card-sorting test. Each card in a deck is defined by two features: color and shape. In each iteration, the agent receives a cue card and two guide cards—one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward $R(t)=1$. Initially, the correct choice is based on color matching. However, after 10 trials, the classification rule changes, rewarding shape matching instead. A successful agent must quickly adapt its decision-making to these contextual shifts.
We consider a simple neural network that predicts reward probability for each (cue card, guide card) pair. Two populations, $z_{1}$ and $z_{2}$, encode similarity based on color and shape, respectively. A third population, $y=\phi(w_{1}z_{1}+w_{2}z_{2})$, integrates this information to predict the reward probability. Here $\phi(x)=\text{sigmoid}(x-\text{bias})$ represents the neural transfer function. The weights $w_{1}$ and $w_{2}$ combine the features to determine the prediction, and the guide card with the highest predicted reward probability is selected. The agent dynamically updates the model based on received rewards to adapt to changing contexts.
Traditionally, weights $w_{1}$ and $w_{2}$ are considered synaptic weights, and learning is attributed to synaptic plasticity. Alternatively, $w_{1}$ and $w_{2}$ can be reinterpreted as the activity of two additional populations acting as gain modulators In our model, we refer to them as "virtual weights," as they replace traditional physical synaptic weights, effectively virtualizing the learning process. These modulatory populations dynamically reconfigure the responses of the network, enabling fast, flexible adaptations without the need for synaptic changes. One of the key advantages of this virtualization is that it allows for learning rules that are not constrained by the strict locality requirements inherent to synaptic plasticity.
To support this point, in Fig. 1 we compare synaptic learning using the gradient-based, delta rule $\Delta\mathbf{w}=\eta e\cdot\mathbf{z}$ with a non-local learning rule that force constraints $w_{1}(t)+w_{2}(t)=const$. This results in the update rule $\Delta w_{j}=\eta(-1)^{j}e\cdot(z_{2}-z_{1})$. Here $e(t)=(R(t)-y(t))$ represents the reward prediction error for the last choice and $\eta$ is the learning rate, controlling the speed of adaptation. As the latter learning rule features non-local terms, it can easily be implemented by a network that dynamically adjusts the activities of two modulatory populations of virtual weights $w_{1},w_{2}$ without relying on synaptic updates (see Fig. 1 C). In Fig. 1 we compare the performance of an agent that learns with synaptic plasticity with an agent that performs dynamic adaptation. As we can see in panel 1 F, decreasing the timescales of the learning process, the network with dynamic adaptation achieves maximum reward after seeing only one in-context example, while the performance of the network with synaptic plasticity plateaus at near chance level. This is supported by the psychometric curve shown in Fig. 1 G. Interestingly, the comparison between the two curves in Fig. 1 G closely mirrors findings from a recent study [30]. In that study, some human participants displayed the ability to adjust their behavior after just one example, while others, with impaired cognitive abilities, had to relearn the previous classification rule through trial and error from scratch. In this work, we propose biologically plausible framework for understanding rapid behavioral adaptations in dynamic environments, bridging the gap between biological and artificial intelligence.
Figure 1: Adapting behavior in a context-dependent classification task. A. Overview of the task structure: At each iteration, the agent is presented with a cue card and two guide cards, one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward. The classification rule switches after 20 trials. B. Schematic representation of the proposed network architecture: synaptic plasticity of physical weights is replaced by dynamic adaptation of virtual weights C. that modulate the network response. D. Example temporal trajectory for the plastic network. Two in-context errors are needed to learn the correct rule. E. Example temporal trajectory for the network with dynamic adaptation: one in-context errors are needed to learn the correct rule. F. Average reward after 1 in-context example as a function of the learning rate $\eta$. Synaptic plasticity with a local delta rule does not go beyond the chance level. Dynamic adaptation with a non-local learning rule can achieve a perfect score. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each. The shaded region represents the standard deviation confidence interval. G. Psychometric curve comparing average reward as a function of in-context examples. At $-1$ we report the average reward before the rule switch. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
Framework Formalization
Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$. The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f$ from the environment in return. This can be an error, reward, or the target behavior.
We further assume that a low-level network extracts features $\mathbf{z}(t)$ relevant to the task family. The agent then adapts its response to the specific environment by learning readout parameters $\mathbf{w}$, which determine how to recombine these features to produce the correct response, $y(t)=Y(\mathbf{z}(t),\mathbf{w})$. Here $Y$ is a generic function that depends on the parameters $\mathbf{w}$. These parameters are then adjusted at each time step in function of an internally computed error signal $e$. The overall learning process can then be generically rewritten as the following dynamical system:
$\begin{cases}\tau_{e}\,\dot{\mathbf{e}}&=E(\mathbf{e},\mathbf{y},\mathbf{f}),\\ \tau_{y}\,\mathbf{y}&=Y(\mathbf{y},\mathbf{w},\mathbf{z}),\\ \tau_{w}\,\dot{\mathbf{w}}&=W(\mathbf{w},\mathbf{z},\mathbf{e}).\end{cases}$ (1)
The specific instantiation of the functions $W$ and $E$ depends on the learning algorithm employed and can encompass both supervised and reinforcement learning settings. Refer to the Methods section for their precise definitions. In the classic view, the parameters $\mathbf{w}$ are interpreted as synaptic weights of a neuronal network, and the functions $E$ and $W$ describe synaptic plasticity mechanisms. However, this imposes stringent locality requirements, as all information must be present at the synapse.
Rather than interpreting $\mathbf{w}$ as synaptic weights, we view them as the dynamic activity of an auxiliary network that modulates the response $Y$ in real time based on feedback and past interactions. We refer to these dynamically adapting state variables as "virtual" weights. We refer to $\mathbf{w}$ as “virtual weights” because: - Weights: they replace the physical weights, and they have the same role as the physical weights of parametrizing the map from the features $\mathbf{z}(t)$ to the output(s). - Virtual: they are technically not weights, since the are encoded in the activity of the readout neurons of the network W, and their activity is provided as an input current to the Y network. Using this strategy we replace the computation done by the synapses (including learning) with operations performed by the neural activity. Dynamic adaptation offers key advantages over synaptic plasticity:
  • Fast Timescales: It naturally operates on faster timescales, making it well-suited for rapid behavioral changes
  • Non-Locality: Dynamic adaptation is not restricted by the stringent locality requirements of synaptic plasticity, allowing for richer forms of interaction and computation at the network level.
In this interpretation, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights. We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning. While these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the circuit level. We demonstrate that such circuits capable of dynamic learning can simply arise by tuning the static readout weights of random RNNs with multiplicative interactions, requiring only a limited number of training trajectories from the original learning algorithm applied to a subset of the task family. The construction of our networks follows two distinct phases: a developmental or evolutionary phase, and a subsequent phase where the network leverages this structure for dynamical learning.
  • Phase 1: Development of the network structure. In this phase, using an Algorithmic Distillation protocol[31], we collect learning trajectories of the form $\{(\mathbf{z}_{\alpha}(t),\mathbf{w}_{\alpha}(t),\mathbf{e}_{\alpha}(t),\mathbf{f}_{\alpha}(t))_{t=0}^{T}\|\alpha\in\mathcal{A}_{ID}\}$ by running the learning dynamics in (1) on a subset of tasks $\mathcal{A}_{ID}$ (In Distribution). We then train the static readout weights $\Theta$ of two random RNNs $Y_{\Theta_{y}}$ and $W_{\Theta_{w}}$ to replicate, respectively, the functions $Y$ and $W$ on the training data.
  • Phase 2: The whole architecture is tested (dynamical learning). The architecture is tested on a new, previously unseen task $\alpha\in\mathcal{A}_{OOD}=\mathcal{A}\setminus\mathcal{A}_{ID}$ (Out-of-Distribution). During this phase, all synaptic weights remain fixed. Instead of relying on synaptic plasticity, the virtual weights evolve dynamically according to: $\mathbf{w}(t+dt)=\mathbf{w}(t)+W_{\Theta_{w}}(t)$. The updated output of the network is given by $y(t)=Y_{\Theta_{y}}(\{\mathbf{w}(t),\mathbf{z}(t)\})$. Importantly, we remark that $\mathbf{w}(t)$ no longer represents physical synaptic weights but rather contextual signals encoded in the activity of neurons, acting as input currents to the $Y_{\Theta_{y}}$ network.
The role of gain-modulation in adaptive behavior
Recent research has highlighted a growing interest in understanding modulatory mechanisms and their role in neural computation. It has been observed that gain modulation allows neurons to adjust their response to inputs based on context [24]. This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility. Moreover, it has been implicated in various functions, including feedback control [32, 33]) and attention, where it allows for the selective amplification of relevant information. A recent study [34] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex. Furthermore, gain modulation is closely tied to meta-learning and multitask learning [35, 36], enabling neural networks to efficiently generalize across diverse tasks by leveraging contextual cues.
Within biological recurrent networks, dendritic segregation provides a structural basis for gain modulation. By compartmentalizing synaptic inputs, neurons can effectively implement non-linear interactions between different input streams. This is reminiscent of In-Context Learning (ICL) in machine learning, where multiplicative interactions are crucial for enabling models to leverage previously learned information in new contexts. Although we focus on gain modulation mediated by apical dendrites, other mechanisms are also known. For instance, neuro-glia interactions at the synapse have been proposed as a substrate for working memory, with astrocytes playing a crucial role in modulating neural activity and maintaining memory through glia–synapse interactions [37]. This gain modulation mechanism could complement other adaptive processes, such as subtle, correlated changes in connectivity [33].
Our work explores the role of gain modulation in recurrent networks, employing a simplified model where apical currents act multiplicatively on basal currents within a transfer function. A key parameter, $\gamma$, controls the strength of these non-linear interactions (see Methods for further details). By modulating the gain of neuronal responses, these networks exhibit enhanced adaptability and in-context learning capabilities.
Results
Figure 2: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of $N=20$ units (described by the vector $\mathbf{z}(t)$, depicted in pink, which follows Eq. (4)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights $\delta w$, a $N$-dimensional vector (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (solid: median, dashed: 20-th/80-th percentile (statistics evaluated over 10 realizations).
Dynamical adaptation for temporal trajectories
We consider the task of autonomously predicting a temporal trajectory $\{f(t)=y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm (see Methods section), we employ a random RNN (Fig. 2A, pink), whose dynamics follows Equation (4), to extract temporal features $\mathbf{z}(t)$. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x(t)=y^{targ}(t)$ in Eq. (4)) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 2B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section. The readout weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},\mathbf{w}\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},e\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\,e\end{cases}$ (2)
[…]
These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$. More specifically, the target sequences are $5$ sinusoidal functions with different frequencies (see vertical red lines in Fig. 2C.)
In the closed-loop phase, the features $\mathbf{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq. (4)). This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 2B). In Fig.2B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line. The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.2C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C, solid and dashed lines represent respectively, median and 20th/80th percentile range. Statistics is evaluated over 10 realizations of the experiment).
Figure 3: Dynamical reinforcement learning: Multi Armed Bandits. A. Schematic of network architecture and task. Virtual weighs parameterize the agent’s policy (orange). At each round, an action (red) is sampled from the policy and played. A reward is then sampled from the current environment. A GM-network (green) then predicts the virtual parameter update. B. Regret comparison between the distilled and the original policy gradient algorithm. We report log regret per round distribution achieved at the 100-th round for 100 independently trained models in ID and OOD settings. We compare a gain-modulated network ($\gamma=1$), and a network without gain modulation ($\gamma=0$) with the policy gradient training source (PG). C. OOD model performance varying number of hidden dimensions. Solid lines indicate the median score (expected reward), computed over 100 independently trained models. The filled area indicates 20-80% confidence interval. D. We report average regret per round curves in OOD setting. We compare a gain-modulated network ($\gamma=1$, teal), and a network without gain modulation ($\gamma=0$, orchid) with the policy gradient training source (in black). Each curve represents the average regret per round for one fixed trained model, computed by averaging over 100 independent simulations.
Extension to Reinforcement learning and the role of Gain modulation
Dynamical adaptation is now investigated within the framework of reinforcement learning. We provide robust evidence supporting the hypothesis that gain modulation represents a fundamental component in implementing in-context learning in biological agents. We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [38]. Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability ($p=0.95$). During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
In this simplified bandit scenario, where state information is absent, the virtual weights $\mathbf{w}$ directly parameterize the policy probabilities: $\mathbf{\pi}=\mathrm{softmax}{(\mathbf{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}^{\gamma}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 3A, green). This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios. The network $W_{\Theta^{w}}^{\gamma}$$(\mathbf{\pi},a,r)$ has a gain-modulated architecture, as described in the Methods section, and is a function of the reward $r$ and the action $a$. Here $\gamma$ indicates the dependence on the gain-modulation strength. We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) in out-of-distribution (OOD) environments. To systematically investigate the impact of gain modulation ($\gamma=1$) on out-of-distribution performance, we compare it with networks where the reward does not modulate the network gain ($\gamma=0$). For each case, we select the optimal hyperparameters through a grid-based search (additional details in the Appendix).
Our experimental findings are presented in Fig.3. First, comparing models of different sizes (number $N$ of hidden units), we find that models with gain modulation require significantly fewer neurons to achieve maximum scores in OOD environments (Fig.3C). In Fig.3B, we report the regret per round distribution at the 100th round and compare it with the distribution obtained by policy gradient with the same learning rate and iterations. A gain-modulated network achieves a regret distribution comparable to the policy gradient in both ID and OOD environments, with a lower median. In contrast, networks without gain modulation show significantly higher regret. Analyzing the regret curves in Fig.3D, we observe that a model with gain modulation often learns a more data-efficient algorithm than its source, even in OOD environments. Conversely, a model without gain modulation fails to generalize to OOD environments and does not converge to zero regret.
In summary, gain modulation enables the network to consistently and efficiently distill the correct gradient update rule and generalize it to unseen environments, predicting the correct virtual weight update in regions of the input space far from the training data.
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). C. Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD). Line: median, shading: 20-th/80-th percentile range. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [39]),
set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\mathbf{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
To accomplish this task we consider a network architecture composed of two networks. One GM-network, $W_{\Theta^{w}}^{\gamma}$($(\{\mathbf{z},r,a,\mathbf{\pi}\})$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 4A, green network), while another GM-network $Y_{\Theta^{y}}^{\Gamma}$($(\{\mathbf{z},\mathbf{w}\})$ estimates policy reconfiguration (Fig. 4A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\Gamma$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equation Eq.(10). Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig. 4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig. 4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions). After multiple trials, the agent successfully achieved precise targeting of the circles. Data collected from these experiments were utilized to train our networks to estimate gradients and scalar products, as defined in Eq. (10).
To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig. 4C, left panel) and new test positions (OOD, Fig. 4C, right panel). For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training. The agent’s precision closely matches that of the plastic policy gradient learning rule. We present the reward plotted against the number of trials for both training (Fig. 4E, right panel, pink line) and test (Fig. 4C, right panel, pink line) food locations.
We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig. 4B, D, F). Notably, while performances for ID food locations are acceptable (Fig. 4E, left panel, pink lines), those for OOD cases are extremely poor (Fig. 4E, right panel, pink lines). This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
Temporal credit assignment and Future Reward
We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section Reinforcement Learning and the Discount Factor in Appendix), requiring evaluating temporal credit assignment. Indeed, in reinforcement Learning (RL) an agent learns to maximize cumulative rewards. The discount factor, usually denoted by $\gamma_{d}$ (where $0\leq\gamma_{d}\leq 1$), is crucial in RL as it determines the importance of future rewards. A higher $\gamma_{d}$ values future rewards more significantly, encouraging the agent to consider the long-term consequences of its actions. Conversely, a lower $\gamma_{d}$ makes the agent prioritize immediate rewards. The cumulative reward $R_{t}$ at time step $t$ is given by:
$R_{t}=\sum_{s\geq t}r_{s}\gamma^{(s-t)}_{d}$ (3)
To account for this, the term $\mathbf{e}$ in Eq.(10) should be replaced by its exponential temporal filtering, with a timescale $\tau_{e}=-\frac{1}{\log(\gamma_{d})}$ (see also Eq.(13) in the Appendix). In this way actions in the past, can be potentiated when delayed rewards are received.
If the discount factor is zero, the agent’s policy would change only when it is very close to the moment the reward is received. In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points (see Fig. A2.D-E). As a result, the policy points towards the target position only when nearby the target itself (see Fig. A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig. A2F, black line).
On the other hand, recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions, achieving optimal planning and decisions even when far from the target location (and hence, from the reward, see Fig. A2.A-B-C).
Discussion
The brain’s ability to flexibly and rapidly adapt to novel tasks remains only partially understood. Different studies have highlighted that this adaptability relies on mechanisms such as dynamic network reconfiguration and context-dependent modulation, allowing agents to respond flexibly to environmental changes. However, a significant gap remains in understanding how these processes map onto specific neural substrates.
While synaptic plasticity has been extensively studied as a mechanism underlying learning and memory [40], emerging evidence suggests that it alone may not account for the full spectrum of adaptive behavior. Abraham et al. critically evaluate the synaptic plasticity and memory (SPM) hypothesis, concluding that non-synaptic mechanisms, such as intrinsic plasticity, structural neural changes, epigenetic regulation, glial contributions, network reorganization, and protein turnover, likely contribute significantly to memory storage and retrieval, thus broadening the framework for understanding dynamic adaptation. This perspective questions long-standing assumptions about neural circuits, highlighting that we may still lack full understanding of many processes we consider well-established.
Inspired by the brain’s adaptability, artificial neural network architectures have been challenged to replicate similar capabilities. Fast synaptic plasticity, hypothesized to orchestrate flexible cognitive functions, inspired several artificial architectures, such as fast-weight networks [41, 42] and hyper-networks that predict main network weights [43]. However, these approaches often rely on biologically unrealistic mechanisms, such as requiring neurons to directly act on the synapses of a second network or implementing "inner loops" in network dynamics [44, 43]. More biologically plausible alternatives, like coupled neural-synaptic dynamics [45], come closer to natural systems but still depend on complex credit assignment mechanisms.
Transformers, known for their remarkable in-context learning (ICL) capabilities, excel at tasks requiring adaptability, such as natural language and image processing [14, 15, 16]. Mechanisms such as attention, associative memory [17], and induction-based copying by attention heads [18] have been proposed to explain ICL in transformers, with theoretical work showing that attention mechanisms can implement gradient updates [19]. However, the emergence of these capabilities remains poorly understood, particularly when comparing them to the in-context learning observed in biological systems.
We propose a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios. We offer an alternative formulation in which dynamic adaptation is modulated by dendritic input segregation and amplification. Virtual (fast) weights are encoded by the activity of a second network that modulates the transfer function of neurons in the base network. As a result, virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other neurons, which were pre-trained to perform gradient updates. Another pre-trained network receives those weights signals as an input, along with the current environment features, and computes an adapted in-context response.
This mechanism overcomes the strict locality constraints of synaptic plasticity, which require all elements involved in evaluating weight updates to be present at the synaptic site. Consequently, biologically plausible synaptic updates cannot explicitly rely on information from other weights, a requirement often seen in backpropagation methods. In contrast, virtual weights are encoded by neuronal activity and can propagate across the neural circuit, bypassing these limitations. This allows virtual weight updates to depend on other virtual weights, enabling a more natural and precise evaluation of credit assignment.
Our work generalizes and provides a biologically plausible implementation for deep gated RNNs performing ICL [46], maintaining a similar number of trainable parameters. By dividing the architecture into two components and introducing an additional objective that forces one network to approximate the gradient of the other, we achieve robust dynamic adaptation in tasks requiring temporal learning. When tasked to learn temporal trajectories, through error feedback and virtual weight updates, our network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, which are updated in function of the reward. We find that networks with gain modulation exhibit improved performance and robustness compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing their capability to face novel scenarios.
While we do not fully address how the static weights of the networks involved can be learned using biologically plausible learning rules, we demonstrate that these architectures naturally emerge by tuning the readout weights of random networks. These weights could be shaped through evolutionary processes, adapted over slow timescales via reinforcement learning, or optimized using local plasticity rules based on feedback alignment. We argue that this represents a step forward in biological plausibility compared to existing deep-learning models like transformers, gated RNNs, and hyper-networks, which rely on more complex and biologically implausible credit assignment mechanisms such as backpropagation through time to train synaptic weights.
In conclusion, our approach provides an explicit framework to induce in-context learning (ICL) in biologically plausible networks, offering a constructive pathway toward understanding ICL in biological systems and bridging the gap between neuroscience and machine learning.
Methods
Gain modulated reservoir computing (GM-RC)
Reservoir Computing
Reservoir Computing (RC) represents a paradigm in machine learning that provides a model for biologically plausible computation and learning, drawing inspiration from the information processing mechanisms observed in biological neural networks.
In this approach, a random RNN is employed to extract features from a time-dependent signal $\mathbf{x}(t)$. The dynamics of the RNN describe the evolution of $N$ hidden units $\mathbf{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
$\tau_{z}\dot{\mathbf{z}}=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}$ (4)
The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([47]). Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively. Subsequently, the features extracted by the network can be utilized to predict a target signal $\mathbf{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\mathbf{y}^{targ}(t)-\Theta\mathbf{z}(t)|^{2}$ is minimized. The reservoir is then implementing a map
$\mathbf{y}(t)=\mathrm{RNN}_{\Theta}(\{\mathbf{x}(s)\}_{s\leq t})$ (5)
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\mathbf{y}=\mathrm{RNN}_{\Theta}(\{\mathbf{x}\})$, dropping the temporal dependencies. 111 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [48].
When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit. In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\mathbf{x}\right)$. Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights. This architecture is also referred to in the literature as the Extreme Learning Machine [49].
Gain modulated network architecture
Analysing Equations (8) and (10) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters. Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC). This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [50, 51]. Here, we consider an additional input source $\mathbf{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$. Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope. Consequently, the resulting RNN equation is formulated as:
$\tau_{z}\dot{\mathbf{z}}$ $=\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(J\mathbf{z}+\beta\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}+R\mathbf{x})\right)-\mathbf{z}$ (6)
Where $\mathbf{b}^{\text{ap}}$ is a constant bias vector. The hyperparameters $\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs. Specifically, when $\gamma=0$, we obtain a network in which $\mathbf{x}^{\text{ap}}$ does not affect the gain modulation. Similarly, as before, the expression $\mathbf{y}=\mathrm{GMRNN}^{\gamma}_{\Theta}(\{\mathbf{x}|\mathbf{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies, we will have an instantaneous function $\mathbf{y}=\mathrm{GMNN}^{\gamma}_{\Theta}(\mathbf{x}|\mathbf{x}^{\text{ap}})=\Theta\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(R\mathbf{x})\right)$ of the inputs $\mathbf{x}^{\text{ap}}$ and $\mathbf{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network. The choice for the name $\mathbf{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [50].
Dynamical supervised learning for temporal trajectory
We consider the task of learning a target temporal trajectory. Each environment $\alpha$ is defined by a target trajectory $y_{\alpha}^{targ}(t)$, such that the input and the feedback $x_{\alpha}(t)=f_{\alpha}(t)=y_{\alpha}^{targ}(t)$ are both given by the target sequence. The network response is defined as the current estimation of $y^{targ}(t)$ via a linear readout $y(t)=\mathbf{w}\cdot\mathbf{z}(t)$, where $\mathbf{z}$ are the activity of a random RNN that operates as a reservoir computer extracting nonlinear features of the input sequence. The RNN dynamics is given by equation (4), with input $x$ given by the target sequence. The readout weights in turn dynamically adapt by following the delta rule, minimizing the reconstruction error.
The learning process can then be described by the following dynamical system:
$\begin{cases}e&=(y^{targ}-y)\\ y&=\mathbf{z}\cdot\mathbf{w}\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{z}\,e\end{cases}$ (7)
where we removed the dependence on time $t$ for simplicity. The operations required are nonlinear, and usually are viewed as implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights. However, here we interpret $\mathbf{w}$ as dynamic variables, proposing, as a possible implementation, that such non-linear functions are computed by two neural networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
$\begin{cases}e&=(y^{targ}-y)\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{x}|\mathbf{w}\})\simeq Y(\mathbf{x},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{x}|e\})\simeq W(\mathbf{x},e)=\mathbf{x}\,e\end{cases}$ (8)
In particular, we used two GMRNNs that receive $\mathbf{x},\mathbf{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm). Notice that now (8) represents an RNN dynamically performing the learning process.
Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behavior across diverse environments. We assume the same state transition dynamics across all the environments $\dot{\mathbf{x}}=X(\mathbf{x},a)$, while the reward distribution $r\sim Prob_{\alpha}(r|\mathbf{x},a)$ depends on the current environment $\alpha\in\mathcal{A}$. Here $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones.
We start by defining a policy network, denoted as $\mathbf{\pi}=\mathrm{softmax}(\mathbf{y})$ which implements a policy mapping the agent state encoded by the feature vector $\mathbf{z}$, to a probability distribution over actions. We assume an agent that linearly recombines these features $\mathbf{y}=Y(\mathbf{z},\mathbf{w})=\mathbf{w}\cdot\mathbf{z}$. This assumption is made without loss of generality, since we could always employ a reservoir computer as an intermediate feature extraction layer, as described in the previous subsection. The agent learns by evaluating the policy gradient with respect to the weights $\mathbf{w}$. This leads to a dynamics similar to equation (8), where the definition of $e(t)$ is changed as follows:
$\begin{cases}\mathbf{e}&=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}\right)\\ \mathbf{y}&=\mathbf{w}\cdot\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{e}\odot\mathbf{z}\end{cases}$ (9)
Here $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [52, 53]) at time $t$. It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero. This formulation holds for a null discount factor [52, 53]), we refer to the Appendix section Reinforcement Learning and the Discount Factor for the description of the general case.
As before, we model that the policy is modulated by an auxiliary population of virtual weights $\mathbf{w}$, determined by the activity of an additional GMRNN $W_{\Theta_{y}}$. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behavior in real-time. Then, the above dynamics can be approximated by two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{z}|r,a,\mathbf{\pi}\})\simeq W(\mathbf{z},r,a,\mathbf{\pi})=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}(\mathbf{y})\right)\odot\mathbf{z}\\ \mathbf{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{z}|\mathbf{w}\})\simeq Y(\mathbf{z},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\end{cases}$ (10)
Acknowledgments
This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’). LF aknowledges support by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing and Sapienza University of Rome (AR12419078A2D6F9). Work partially funded by the European Union – NextGenerationEU and the Ministry of University and Research (MUR), Italian National Recovery and Resilience Plan (PNRR), M4C2I1.3, project ‘MNESYS’ (PE00000006). We sincerely thank Maurizio Mattia for helpful discussions on the topic.
Source code availability
The source code is available under CC-BY license in the https://github.com/cristianocapone/ABSS public repository.